OpenStudy (anonymous):

help!?! A snow cone is a tasty treat with flavored ice and a spherical bubble gum ball at the bottom.

OpenStudy (anonymous):

The radius of the cone is 1.5 inches, and its height is 3 inches. If the diameter of the bubble gum ball is 0.5 inches, what is the closest approximation of the volume of the cone that can be filled with flavored ice? 21.21 in3 0.07 in3 7.00 in3 7.07 in3

OpenStudy (anonymous):

@travisbrown372 help plz

OpenStudy (anonymous):

@esshotwired help plz

OpenStudy (anonymous):

@campbell_st help plz

OpenStudy (campbell_st):

well find the volume of the cone \[V = \frac{1}{3} \pi r^2 h\] and then subtract the volume of the sphere \[V = \frac{4}{3} \pi r^3\] you have been given all the necessary measurements so substitute to find the volumes then subtract for the approximation.

OpenStudy (esshotwired):

plug in the numbers to the equations that campbell_st gave you above, then i will tell if you got it right or not

OpenStudy (anonymous):

@esshotwired is this right